OpenStudy (anonymous):

Find dy/dx .... xy^2 = cos x

OpenStudy (usukidoll):

implicit differentiation?!

hartnn (hartnn):

implicit differentiation do you know product and chain rule ?

OpenStudy (anonymous):

Is that like d/dx (xy^2) = d/dx cos x

OpenStudy (anonymous):

Product rule: \[\huge f'(x)g(x) +f(x)g'(x)\]

OpenStudy (anonymous):

\[(x \frac{ d }{ dx } (y ^{2}) + y ^{2} \frac{ d }{ dx } (x)\]

OpenStudy (anonymous):

Chain rule, here to make it more clear: \[\huge h(x) = f[g(x)]\] \[\huge h’(x) = f’[g(x)]g’(x)\]

OpenStudy (anonymous):

but because there is a y what do you do?

OpenStudy (anonymous):

dy/dx

OpenStudy (usukidoll):

don't you have to solve for dy/dx because in the end dy/dx will always be on the left hand side of the equation..

OpenStudy (anonymous):

Yes, that's the point :p

OpenStudy (usukidoll):

I would just solve out right but nahhhhhhhhh where is the learning ? :P

OpenStudy (anonymous):

If it's easier, just use y' notation.

OpenStudy (usukidoll):

NO! that's BAD S:

OpenStudy (usukidoll):

who wants me to just end the misery type 1

OpenStudy (anonymous):

wait...im still confused

OpenStudy (usukidoll):

hint: every time you take the derivative of y you have to write dy/dx next to it

OpenStudy (anonymous):

What do you get when you use product rule on xy?

OpenStudy (usukidoll):

type 1 to end misery ^^.

OpenStudy (anonymous):

Yeah pretty much what usuki said, dy/dx wherever you're finding derivative of y.

OpenStudy (usukidoll):

another hint: you need the chain rule and product rule... if you haven't mastered it yet go practice then come back to this

OpenStudy (usukidoll):

for example y^2 derivative of y 2y dy/dx

OpenStudy (anonymous):

The rules are also written above ^^

OpenStudy (anonymous):

\[xy ^{2} = \cos x\] \[\frac{ d }{ dx } (xy ^{2}) = \frac{ d }{ dx } (\cos x)\] \[(x \frac{ d }{ dx } (y ^{2}) + y ^{2} \frac{ d }{ dx } (x) = -\sin x\]

OpenStudy (usukidoll):